A study was conducted in a large population of adults concerning eyeglasses for correcting reading vision. Based on an examination by a qualified professional, the individuals were judged as to whether or not they needed to wear glasses for reading. In addition it was determined whether or not they were currently using glasses for reading.The following table provides the proportions found in the study.

a. If a single adult is selected at random from this large population, what is the probability that the adult is judged to need eyeglasses for reading?

b. What is the probability that a selected adult is judged to need eyeglasses but does not use them for reading?

c. Suppose two adults are selected from the population independently and at random. What is the probability thatboth were judged to need eyeglasses and neither was using them for reading?

Used glasses for reading

Yes

No

Judged to need glasses

Yes

0.42

0.18

No

0.04

0.36
Answer #1

a)

P( adult is judged to need eyeglasses for reading ) =0.42+0.18=0.60

b)

probability that a selected adult is judged to need eyeglasses but does not use them for reading =0.18

c)

probability that both were judged to need eyeglasses and neither was using them for reading =0.18*0.18=0.0324